OpenStudy (anonymous):

find the derivative of x^2/root x-4

OpenStudy (calculusfunctions):

\[\frac{ d }{ dx }(\frac{ x ^{2} }{ \sqrt{x -4)} })=?\]This is your question, correct?

OpenStudy (anonymous):

yes

OpenStudy (calculusfunctions):

OK then first rewrite as\[y =\frac{ x ^{2} }{ (x -4)^{\frac{ 1 }{ 2 }} }\]and use the quotient rule, or rewrite as\[y =x ^{2}(x -4)^{-\frac{ 1 }{ 2 }}\]and use the product rule. Which rule is your preference?

OpenStudy (anonymous):

i used the quotient rule and got (root x-4) (2x)-(x^2)-(1/2root x-4)/ (root x-4)^2. And now I don't know what to do from here... I don't know if i'm using the quotient right or not.

OpenStudy (calculusfunctions):

Which means so far you got\[\frac{ dy }{ dx }=\frac{ 2x \sqrt{x -4}+\frac{ x ^{2} }{ 2\sqrt{x -4} } }{ x -4 }\]Understand?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait isn't it suppose to be a minus in the middle of the numerator instead of the add? Why are you adding?

OpenStudy (calculusfunctions):

negative times a negative, or a negative divided by a negative equals positive. Right?

OpenStudy (anonymous):

whoops, made a mistake my previous answer is suppose to be (root x-4) (2x)-(x^2)(1/2root x-4)/(root x-4)^2. There is technically no negatives.

OpenStudy (anonymous):

only the minus sign due to the quotient rule.

OpenStudy (calculusfunctions):

Oh yes you're right actually, it is a minus sign.

OpenStudy (calculusfunctions):

\[\frac{ dy }{ dx }=\frac{ 2x \sqrt{x -4}-\frac{ x ^{2} }{ 2\sqrt{x -4} } }{ x -4 }\]

OpenStudy (calculusfunctions):

OK now you need to simplify this. Do you know how?

OpenStudy (anonymous):

no...that's where i'm having trouble at.

OpenStudy (anonymous):

multiply top and bottom by \(2\sqrt{x-4}\) and it will simplify like mad

OpenStudy (calculusfunctions):

As satellite said, multiply both the numerator and the denominator by\[2\sqrt{x -4}\]

OpenStudy (anonymous):

is it 4(x-4)-x^2/x(x-4)^3/2?

OpenStudy (calculusfunctions):

No, it's\[\frac{ dy }{ dx }=\frac{ 4x(x -4)-x ^{2} }{ 2\sqrt{(x -4)^{3}} }\]

OpenStudy (calculusfunctions):

The exponent to the root in the denominator is a 3. The equation feature makes it difficult to distinguish between exponents 2 and 3. Alright?

OpenStudy (anonymous):

okay, now what?

OpenStudy (calculusfunctions):

Simplify the numerator. Apply the distributive property and collect like terms.

OpenStudy (calculusfunctions):

\[\frac{ dy }{ dx }\frac{ 4x ^{2}-16x -8x ^{2} }{ 2\sqrt{(x -4)^{3}} }\]

OpenStudy (calculusfunctions):

Now please tell me that you at least know how to collect like terms in the numerator.

OpenStudy (anonymous):

yes I do. Thank you so so much for your help!

OpenStudy (calculusfunctions):

There should be an equal sign there. I must have inadvertently missed it.

OpenStudy (calculusfunctions):

Welcome!